OpenStudy (anonymous):

[2.07] Create an absolute value equation that has two solutions. Explain why this equation has two solutions.

OpenStudy (sw050399):

Well, this is kind of just like the last question you asked. For an easy equation, just change the values of the last problem you posted. Then, just solve it as some users explained on your last question. Let me know what you come up with or if you need a clearer explanation :)

OpenStudy (anonymous):

can you give me a clreare explanation this isn't a multiple answert question and i want to get it right

OpenStudy (agent0smith):

Almost any absolute value equation will have two solutions. eg. \[\large \left| x \right| = a\] \[\large x = \pm a\] Just replace a with any number.

OpenStudy (agent0smith):

Because an absolute value sign changes a negative number into a positive.

OpenStudy (anonymous):

ok so l2l=a ?

OpenStudy (sw050399):

Yep.

OpenStudy (agent0smith):

Yeah i guess that works too... can you find the two solutions for a?

OpenStudy (anonymous):

i dont know how im trying to understand through what hes taught me but im taking the pretest now i havent learned this yet

OpenStudy (sw050399):

Absolute value works in a way that accounts for both the positive and negative. So, whenever you work with absolute value, you just have to remember that you're kinda solving 2 different equations. Example: |x + 5| = 2 You'd get x + 5 = 2 and x + 5 = -2 in order to account for the absolute value.

OpenStudy (anonymous):

thats why you get two solutions ? to accout for the absolute value ?

OpenStudy (sw050399):

Yep. Of course there are some weird times where it will be the 2 answers will be the same, so you just need to split it up into 2 separate equations as I did above.

OpenStudy (anonymous):

oh ok ok im kinda understanding it now